OpenStudy (anonymous):

solve the system by the substitution method 27x-3y=27 y=9x+9

OpenStudy (anonymous):

27x - 3(9x +9) = 27 27x - 27x -27 = 27. 0x -27 = 27 0=54. No solution? Or you typed it incorrectly.

OpenStudy (anonymous):

27x-3y=27 y=9x+9 27x-3(9x+9)=27 27x-27x-27=27 -27=27 => -27 is not the same as 27